OpenStudy (anonymous):

Use the matrix tool to solve the system of equations. Choose the correct ordered pair. 4x + y = 0 8x - y = 6 A. (2, -8) B. (1, -4) C. (,1/2 -4) D. (,1/2 -2)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

any other answers I wanna make sure this is corrcet

OpenStudy (tommo_lcfc):

The inverse of a 2 by 2 matrix [a b; c d] is [d -b; -c a]. So switching the elements on the main diagonal and inverting the signs of the other two gives the inverse as [-1 -1; -8 4]. The determinant of a 2 by 2 matrix is ad-bc = (4)(-1) - (8)(1) = -12. So the inverse is 1/-12 [-1, -1; -8, 4]. Multiplying this by the solution matrix [0,6] gives 1/-12 [6,24] so x = 1/2 and y = -2.